# Large Language Models and Games: A Survey and Roadmap

### v1

@@ 1 Introduction
Five years ago, autoregressive language modeling was a somewhat niche topic within natural language processing. Training models to simply predict text based on existing text was considered something of primarily theoretical interest, though it might have applications as writing support systems. This changed drastically in 2019 when the GPT-2 model was released [1]. GPT-2 demonstrated convincingly that transformer models trained on large text corpora could not only generate surprisingly high-quality and coherent text, but also that the text generation could be controlled by carefully prompting the model. Subsequent developments, including larger models, instruction fine-tuning, reinforcement learning from human feedback [2], and the combination of these features in ChatGPT in late 2022, turbocharged interest in large language models (LLMs). It suddenly appeared that LLMs could do almost anything—as long as both problem and solution could be formulated as text.
LLMs are currently a very active research field, where researchers are both focused on improving the capabilities of LLMs while reducing their compute and memory footprint, and on understanding and learning to harness the capabilities of existing LLMs. Informed opinions on the ultimate capabilities of LLM technology vary widely, from those who see such models as “sparks of AGI” [3] to those who see them as largely doing approximate retrieval from a lossy compression of the internet [4].
Games, including board games and video games, serve both as a source of important benchmarks for AI research and as an important application area for AI techniques [5]. Almost any game utilizes some kind of AI technology, and we are currently in an exploratory phase where the developers and researchers try to figure out how to best make use of recent advances in AI [6]. One might also make the argument that video game design and video game technology might underlie much of the development of future human-computer interfaces and, for lack of a better word, the metaverse.
@@ 2 A Note on Terminology
This paper concerns the intersection between games, including board games, video games, and other types of games people play, and large language models. But what exactly is an LLM?
@@
Broadly speaking, an LLM is a model that is trained on text so as to be able to reproduce text in response to other text. But this definition is overly broad, as it would include Shannon’s original $n$-gram models from 1946 [8], rudimentary recurrent neural nets from the early 1990s [9], and the Tegic T9 text prediction system that would help you write text messages with your Nokia 3210. For readers who are unfamiliar with such technologies, they were ubiquitous in 90’s mobile phones.
@@
What distinguishes LLMs from other text generative models is that they are large. But which model size is considered large enough? LLMs became a well-recognized term with the introduction of OpenAI’s GPT-2 network, released in 2019 [1]. GPT stands for Generative Pretrained Transformer, where transformer is a type of neural network introduced in 2017 [10]. This model became very influential because of what was perceived as a quantum leap in output quality compared to previous models. The various versions of GPT-2 have between 117 millions and 1.5 billions of parameters. Because of the association between the term LLM and the GPT-class of models, we will use the size of GPT-2 as a soft cutoff on the type of models we consider as LLMs; we are concerned with models of few hundred million parameters or more. While language models could in principle be based on various architectures, including LSTM networks, the current LLM landscape is dominated by variants of the transformer architecture; in this survey we rely primarily on this type of model architecture.
@@
It is important to note that LLMs are by no means limited to the GPT family of models. There is by now a large variety of LLMs of varying size and capabilities, including open-source models such as Mistral [11] and the Llama [12] family, which can be fine-tuned and analyzed in various ways and run locally, and even be embedded in games’ runtimes.
@@
One could also argue that the definition is somewhat narrow, as many modern LLMs are multimodal models, meaning that they can take as input and/or produce as output modalities other than text. In particular, many modern LLMs can process and produce images. This is often achieved through combining the core transformer network with a convolutional network for input and a diffusion model for output. Examples include GPT-4V [13] and the open-source Llava [14]. In this paper, we consider large multimodal models (LMMs) [13] as long as they retain their ability to both consume and produce text.
@@
This survey will not concern itself with AI and machine learning techniques that are not LLMs as defined above. In particular, we will not be covering the large literature on game playing and content generation using machine learning methods [15] that is not using textual input and output. We will, however, occasionally mention some of that work when relevant, in particular to help provide historical context.
@@ 3 Roles of LLMs in Games
Past attempts at a typology for AI in games focused on three roles the AI can take in a game: to play a game, to design a game, or to model the (human) players [5]. LLMs are typically presented as conversational agents, which often lead the public to give them anthropomorphic qualities—such as reasoning and creativity. Therefore, we follow these trends when considering the roles an LLM can be called to play within the game or within the game development process. An LLM can operate within the game as a player (replacing a human player while imitating their goals), as a non-player character such as an enemy or interlocutor, as an assistant providing hints or handling menial tasks for a human player, as a Game Master controlling the flow of the game, or hidden within the games’ ruleset (controlling a minor or major mechanic of the game). There are however other roles an LLM can play outside of the game’s runtime, such as a designer for the game (replacing a human designer) or as an assistant to a human designer. Finally, the LLM can interface with a player or an audience in different ways, acting as a commentator of an ongoing play session (during runtime) or a reteller of past game events in some narrative form (outside runtime). Some of these roles (autonomous player, autonomous designer) are prominent in the broader AI and games research [5] and LLM research has targeted them extensively, while some of the other roles have been toyed with in exploratory research. The following sections present the roles themselves, surveying research undertaken for each role, while we identify gaps and opportunities for future research in Section 4.
@@ 3.1 Player
How can an LLM play a game? Fundamentally, language model players require some transformation from their typical output space (i.e. sequences of tokens) into the input space of the game. In addition, aspects of the game and its current state must be provided to the LLM in some form in order for it to play at any kind of reasonable level. Depending on the game itself, these mappings might be intuitive or complex. We identify three general classes of games to which LLM players are well suited: (a) games where states and actions can be compactly represented as sequences of abstract tokens, (b) games where the main input and output modalities are natural language, and (c) games for which external programs can control player actions via an API.
@@
The first class of games mostly includes turn-based board games (e.g. Chess), since the discrete set of board positions and moves is more easily transformed into a compact representation (e.g. Portable Game Notation) than, for instance, a first-person shooter. By tokenizing sequences of moves taken from a game database, the problem of action selection can then be mapped to the standard autoregressive learning objective on which LLMs are trained – predicting the next move given the context of those that preceded it. Chess [16, 17, 18], Go [19], and Othello [20] have all been used as testbeds for LLM players in this way. However, board games are not the only kind of game that can be represented as token sequences: the generalist GATO [21] agent can play a variety of Atari games at human or near-human levels by processing visual inputs as sequences of pixel values in raster order. Pixel values are interleaved with separator tokens and previous actions, allowing the model to accurately predict the appropriate game action in a dataset of human play traces. It is possible that continued improvement in Transformer models that capture both spatial and visual dynamics [22, 23] could allow for a similar approach to scale to even more complex games. However, such approaches require large datasets of gameplay video that may be comparatively more difficult to collect. In addition, we note that the reliance on human gameplay traces as the basis for learning may make it more difficult for an LLM player to reach super-human performance without leaps in terms of reasoning and generalization, which we revisit in the limitations of LLMs (see Section 5).
@@
The second class of games most obviously includes text adventure games such as Zork (Infocom, 1977), where game states are presented as natural language descriptions and the game is already equipped with a parser to handle natural language responses. This means that LLMs can be queried for game actions in a way that still leverages their large-scale pretraining on natural language text. The earliest application of LLMs to these kinds of text games is CALM [24], a GPT-2 system finetuned on a dataset of human gameplay transcripts collected from a variety of text-based adventure games. The model is trained to predict the natural language string provided by human players given the context of previous states and actions and information about the avatar (e.g. their inventory). To actually play a game, the trained language model generates multiple candidate actions and deep reinforcement learning (RL) is used to optimize a policy that selects actions from among the candidates. At the time of its publication, this RL component was necessary because the LLM alone was not capable of generalizing well to unseen games or situations [24]. However, a more recent investigation of ChatGPT as a Zork player has indicated that LLM performance is increasing [25]. In a preliminary experiment, Yao et al. [24] show that the performance of ChatGPT can approach that of existing algorithms for text game playing, as long as a human interlocutor remains in the loop to assist the model (e.g. by reminding it of actions it has already tried). However, there is obviously much room for improvement in directly applying LLMs to text games in this way. Additionally, the ability for LLMs to play entirely novel, niche, or unseen text games (especially important given the likelihood that such systems encounter walkthroughs or playtraces of popular text games during their training) remains largely unexplored.
@@
Beyond text adventure games, the most well-known application of LLM players in board games is CICERO [26] for playing the deal-making and subterfuge game Diplomacy (Avalon Hill Games, 1976). Diplomacy is well-positioned for natural language instruction, as players can freely converse in private messages in order to arrange plans, negotiate, or deceive. CICERO builds from a pre-trained LLM and is fine-tuned on a large corpus of Diplomacy transcripts. Throughout the game, samples from the model are sent to other players and the various dialogue transcripts are collected to condition the potential action. CICERO is further trained to condition its outputs on specific game intents (inferred from the transcripts and added as additional context during training). In order to select an action, CICERO uses as “strategic reasoning module” that predicts the actions of other players, using a value and policy function learned from self-play. Diplomacy is an interesting game in part because the action space is split between natural language utterances and a more standard set of moves on a discrete game board, and CICERO demonstrates how an LLM can be integrated as part of a larger system for high-level play.
@@
Finally, we consider games for which a robust API exists. This is less of a kind of game in the sense of its style or mechanics, and more a fact about its popularity or its ease of implementation. An API is an important attribute because it allows LLMs to act as players not by directly generating actions, but by producing programs that act as policies. Improvements in the code-generation abilities of LLMs have allowed them to write small programs that can produce actions given game states without further intervention from the model. For instance, the VOYAGER system [27] leverages the code-generation abilities of GPT-4 to play Minecraft (Mojang Studios, 2011) by interacting with the popular Mineflayer API. Using a sophisticated chain of prompts, VOYAGER generates blocks of code that leverage calls to the API in order to execute high-level “skills” (e.g. “Attack nearest zombie”) that are automatically converted into low-level game inputs (e.g. mouse movements and key presses). GPT-4 is also used as a high-level goal generator and planner, which in turn informs the code generation. This approach has proved very successful, with VOYAGER being the first automated system to complete a variety of in-game Minecraft challenges. The results are impressive and indicate that generating action-producing programs may be a more efficient way to leverage latent LLM knowledge than direct action sampling. However, VOYAGER does benefit substantially from the availability of a robust API and vast amounts of internet discussion for its target game of Minecraft. As with the analysis of ChatGPT on Zork, the ability of this approach to generalize to less popular or entirely unseen games remains to be seen.
@@ 3.2 Non-Player Characters
Non-player characters (NPCs) are agents which exist in virtual game worlds but whose actions are not directly controlled by the players. NPCs exist to enrich the player’s experience and deepen immersion by adding to the world’s ambiance and making it more believable [29]. NPCs may serve as pets, allies, enemies [^1] We use enemies and allies here for in-game agents with different skills and ways to affect the world than e.g. gameplaying AI such as opponents in Chess., merchants, quest givers, or bystanders. Therefore, they have different agency even from AI-controlled players, and their goal is never to win. This makes designing AI for NPCs interesting [5], while LLMs can provide unique advantages in this task. Their ability to “understand” game-world settings allows LLMs to adapt their responses to match these settings. It has been shown that LLMs are able to role-play through different scenarios [30], thereby highlighting their potential to provide a more flexible and apt tool to emulate human behavior. We identify two ways in which LLMs can control NPCs: (a) through their dialogue, and (b) through their behavior. Behavior relates to in-game action selection, discussed in Section 3.1; however, we note that the heuristics and goals of such behavior is different than an AI player trying to win the game.
@@
LLMs are naturally suited for natural language conversation, and as NPC dialogue systems they can generate dynamic and contextually appropriate responses based on player input. This makes interactions with NPCs more engaging and realistic, reduces repetitive discourse and provides a more explorative experience within the game [31]. LLMs can engage the players in the gameworld’s narrative as foreground NPCs, background NPCs, or narrator NPCs. We discuss narrator LLMs as commentators in Section 3.4 whereas we cover the other two NPC types here. Foreground NPCs form part of the overarching narrative of the game, or one of its sub-narratives. They may be enemies, allies, information-givers, quest-givers, or item-providers. Their dialogue is heavily constrained by the scope of the narrative, their role within it, and the player actions. Foreground NPCs’ text generation process via LLMs must consider the overall context of the game and the interaction with the player and keep track of events transpiring in the playthrough. This raises concerns regarding the memory capacity of LLMs, as well as the impact of possible hallucinations, i.e. plausible but false statements [32]; we revisit these limitations in Section 5.
Background NPCs populate the majority of the virtual world but are not part of any specific task the player is carrying out. Their purpose is to make the environment more believable and act independently of the players [31]. Since such NPCs’ presence is purely decorative and their dialogue is essentially small talk, their dialogue generation is less constrained, perhaps bound only by the identity of the speaker and their background. That said, their believability hinges on their ability to maintain the illusion that they have their own agency in the world and can interact with it [33].
@@
Studies have shown that multiple agents are able to follow game rules and engage in game playing [34, 35], with different models consistently exhibiting their own aptitudes and weaknesses when applied to specific roles. This ability to interact within constraints is useful to instill believable behaviors in foreground and background NPCs, grounding their actions and dialogue within the rules of the game environment. Other work has focused more on the conversational and story-writing abilities of LLMs, such as the creation of dialogue between multiple characters, each having their unique personality, whilst following a consistent plot. One such example is the use of LLMs to generate a South Park (Comedy Central, 1997) episode [36] with multiple characters within a well-known setting. There are limitations to this approach, primarily that LLMs perform something like a theatrical improvisation, rather than acting as an actor studying a part [30]. Through this unconstrained process, the LLM is prone to hallucinations which do not fit the desired scenario. This volatility can be mitigated by providing the LLM not only with the conversation history but also with the current state of the environment, such as the items within it and their affordances, as well as the other characters and their corresponding actions. This method has been successfully employed to create conversational NPCs in text-based games [37], but may also be extended to other scenarios or to cover the use of LLMs as active or interactive narrators.

### v2

@@ 1 Introduction
Five years ago, autoregressive language modeling was a somewhat niche topic within natural language processing. Training models to simply predict text based on existing text was considered something of primarily theoretical interest, although it might have applications as writing support systems. This changed drastically in 2019 when the Generative Pre-trained Transformer 2 (GPT-2) model was released [1]. GPT-2 demonstrated convincingly that transformer models trained on large text corpora could not only generate surprisingly high-quality and coherent text, but also that text generation could be controlled by carefully prompting the model. While not the first autoregressive model [2, 3], GPT-2 was the first of the “large” models, and as such we use it here as cutoff mark (see also Section 2). Subsequent developments, including larger models, instruction fine-tuning, reinforcement learning from human feedback [4], and the combination of these features in ChatGPT in late 2022, turbocharged interest in large language models (LLMs). Capabilities of LLMs were seemingly unbounded —as long as both problem and solution could be formulated as text.
LLMs are currently a very active research field. Researchers are focused on improving the capabilities of LLMs while reducing their compute and memory footprint, but also on understanding and learning to harness the capabilities of existing LLMs. Informed opinions on the ultimate capabilities of LLM technology vary widely, from the enthusiastic [5] to the pessimistic [6]. Our aim is to approach the topic from somewhere in-between these two perspectives: optimistic with respect to the potential of LLMs and realistic with respect to their technical, theoretical, and ethical shortcomings.
Games, including board games and video games, serve both as a source of important benchmarks for research in Artificial Intelligence (AI) and as an important application area for AI techniques [7]. Almost every game utilizes some kind of AI technology, and we are currently in an exploratory phase where both developers and researchers try to figure out how to best make use of recent advances in this field [8].
@@
It is important to note that this survey emerges from the top down, based on our expertise in AI and games [7], and extensive work on most topics covered by this paper. The focus of the paper, in Section 3, is built from our own typology and supported where possible by academic and non-academic work. While a bottom-up approach via e.g. keyword search through general paper repositories is valuable [10, 11, 12], the process would lead to a very different type of paper. However, we have attempted to conduct a comprehensive manual review of all recent proceedings from the major conferences in AI and games [^1] IEEE Conference on Games, Foundations of Digital Games conference, Artificial Intelligence and Interactive Digital Entertainment conference., and the IEEE Transactions on Games for work relevant to the themes of this paper.
@@ 2 A Note on Terminology
This paper concerns the intersection between games (board games, video games, or other), and large language models. But what exactly is an LLM?
@@
Broadly speaking, an LLM is a model that is trained on text in order to be able to reproduce text in response to other text. But this definition is overly broad, as it would include Shannon’s original $n$-gram models from 1946 [13], rudimentary recurrent neural nets from the early 1990s [14], and the Tegic T9 text prediction system that would help you write text messages on your Nokia 3210.
@@
What distinguishes LLMs from other text generative models is mainly that they are large. But which model size is considered large enough? In 2019 emerging models such as BERT [2] and ERNIE [3] showcased significant advances in language modeling, but LLMs became a well-recognized term with the introduction of OpenAI’s GPT-2  [1], whose various versions have between 117 millions and 1.5 billions of parameters. Because of the association between the term LLM and the GPT-class of models, we will use the size of GPT-2 as a soft cutoff on the type of models we consider LLMs; we are concerned with models of few hundred million parameters or more. Each subsequent iteration of the GPT family features an increased number of parameters, and larger and more diverse training corpora.
Another distinct trait of LLMs is their architecture. While language models could in principle be based on various architectures, including Long Short-Term Memory (LSTM) networks [15], the current LLM landscape is dominated by variants of the transformer architecture, a type of neural network introduced in 2017 [16]. This model became very influential because of what was perceived as a quantum leap in output quality compared to previous models. In this survey we rely primarily on LLMs employing this architectural basis.
The last feature of LLMs we consider is their versatility across a wide range of tasks with minimal or no fine-tuning or retraining. This capability represents a significant shift; since the release of GPT-3.5, LLMs have evolved from primarily autoregressive predictive text models to pre-trained, general-purpose conversational models.
@@
It is important to note that LLMs are by no means limited to the GPT family of models. There is by now a large variety of LLMs of varying size and capabilities, including open-source models such as Mistral [17] and the Llama [18] family, which can be fine-tuned, run locally, and even be embedded in games’ runtimes.
@@
One could also argue that the definition is somewhat narrow, as many modern LLMs are multimodal models, meaning that they can take as input and/or produce as output modalities other than text. In particular, many modern LLMs can process and produce images. This is often achieved through combining the core transformer network with a visual encoder network for input and a latent diffusion model for output. Examples include GPT-4V [19] and the open-source Llava [20]. In this paper, we consider large multimodal models (LMMs) [19] as long as they retain their ability to both consume and produce text.
@@
This survey will not concern itself with AI and machine learning techniques that are not LLMs as defined above. In particular, we will not be covering the large literature on game playing and content generation using machine learning methods [21] that does not use textual input and output. We will, however, occasionally mention some of that work where relevant, in particular to help provide historical context.
@@ 3 Roles of LLMs in Games
Past attempts at a typology for AI in games focused on three roles the AI can take in a game: to play a game, to design a game, or to model the (human) players [7]. LLMs are typically presented as conversational agents, which often invites the public to give them anthropomorphic qualities —such as reasoning and creativity. We follow these trends when considering the roles an LLM can be called to play within the game or within the game development process. An LLM can operate within the game as a player (replacing a human player while imitating their goals), as a non-player character such as an enemy or interlocutor, as an assistant providing hints or handling menial tasks for a human player, as a Game Master controlling the flow of the game, or hidden within the games’ ruleset (controlling a mechanic of the game). There are however other roles an LLM can play outside of the game’s runtime, such as a designer for the game (replacing or assisting a human designer) or as analyst of the gameplay data of the player base. Finally, the LLM can interface with a player or an audience in different ways, acting as a commentator of an ongoing play session (during runtime) or a reteller of past game events in some narrative form (outside runtime). Some of these roles (autonomous player, autonomous designer) are prominent in the broader AI and games research [7] and LLM research has targeted them extensively, while some of the other roles have been toyed with in exploratory research. The following sections present the roles themselves, surveying research undertaken for each role, while we identify gaps and opportunities for future research in Section 4.
@@ 3.1 Player
How can an LLM play a game? Fundamentally, LLM players require some transformation from their typical output space (i.e. sequences of tokens) into the input space of the game. In addition, aspects of the game and its current state must be provided to the LLM in some form in order for it to play at a reasonable level. Depending on the game itself, these mappings might be intuitive or complex. We identify three general classes of games to which LLM players are well suited: (a) games where states and actions can be compactly represented as sequences of abstract tokens, (b) games where the main input and output modalities are natural language, and (c) games for which external programs can control player actions via an API.
@@
The first class of games mostly includes turn-based board games (e.g. Chess), since the discrete set of board positions and moves is more easily transformed into a compact representation (e.g. Portable Game Notation [22]) than, for instance, a first-person shooter. By tokenizing sequences of moves taken from a game database, the problem of action selection can then be mapped to the standard autoregressive learning objective on which LLMs are trained —predicting the next move given the context of those that preceded it. Chess [23, 24, 25], Go [26], and Othello [27] have all been used as testbeds for LLM players in this way. This approach allows even more complex game states to be reasoned upon by an LLM player. In Bateni and Whitehead’s work [28], for example, the LLM plays the popular video game Slay the Spire (Mega Crit, 2017), understanding synergies between cards based solely on their description and adapting to changes in gameplay rules. However, board games are not the only kind of game that can be represented as token sequences: the generalist GATO [29] agent can play a variety of Atari games at human or near-human levels by processing visual inputs as sequences of pixel values in raster order. Pixel values are interleaved with separator tokens and previous actions, allowing the model to accurately predict the appropriate game action in a dataset of human play traces. It is possible that continued improvement in transformer models that capture both spatial and visual dynamics [30, 31] could allow for a similar approach to scale to even more complex games. However, such approaches require large datasets of gameplay videos that may be comparatively more difficult to collect. In addition, we note that the reliance on human gameplay traces as the basis for learning may make it more difficult for an LLM player to reach super-human performance without leaps in terms of reasoning and generalization (see Section 5).
@@
The second class of games most obviously includes text adventure games such as Zork (Infocom, 1977), where game states are presented as natural language descriptions and the game is already equipped with a parser to handle natural language responses. This means that LLMs can be queried for game actions in a way that still leverages their large-scale pre-training on natural language text. The earliest application of LLMs to these kinds of text games is CALM [32], a GPT-2 system finetuned on a dataset of human gameplay transcripts collected from a variety of text-based adventure games. The model is trained to predict the natural language string provided by human players given the context of previous states, actions and information about the avatar (e.g. their inventory). To actually play a game, the trained language model generates multiple candidate actions and deep reinforcement learning (RL) is used to optimize a policy that selects actions from among the candidates. At the time of its publication, this RL component was necessary because the LLM alone was not capable of generalizing well to unseen games or situations [32]. However, a more recent investigation of ChatGPT as a Zork player has indicated that LLM performance is improving [33]. In a preliminary experiment, Yao et al. [32] show that the performance of ChatGPT can approach that of existing algorithms for text game playing, as long as a human interlocutor remains in the loop to assist the model (e.g. by reminding it of actions it has already tried). However, there is obviously much room for improvement in directly applying LLMs to text games in this way. Additionally, the ability for LLMs to play entirely novel, niche, or unseen text games (especially important given the likelihood that such systems encounter walkthroughs or playtraces of popular text games during their training) remains largely unexplored.
In a similar vein, inductive biases from large language models can be applied to help guide the policies of agents trained with other methods. For instance, the GALAD system [34] uses a pre-trained LLM to guide an agent towards morally acceptable actions in text games from the Jericho suite [35], while the MOTIF system [36] learns an intrinsic reward function for NetHack NetHack, 1987) by mining preferences between game states from an LLM.
@@
Text adventure games are not the only cases where natural language input and output are used for playing: many boardgames operate via player negotiation. CICERO [37] leveraged LLMs for playing the deal-making and subterfuge game Diplomacy (Avalon Hill Games, 1976). CICERO builds from a pre-trained LLM and is fine-tuned on a large corpus of Diplomacy transcripts. Throughout the game, samples from the model are sent to other players and the various dialogue transcripts are collected to condition the potential action. CICERO is further trained to condition its outputs on specific game intents (inferred from the transcripts and added as additional context during training). In order to select an action, CICERO uses a “strategic reasoning module” that predicts the actions of other players, using a value and policy function learned from self-play. Diplomacy is an interesting game in part because the action space is split between natural language utterances and a more standard set of moves on a discrete game board, and CICERO demonstrates how an LLM can be integrated as part of a larger system for high-level play.
@@
Finally, we consider games for which a robust API exists. This is less of a kind of game in the sense of its style or mechanics, and more a fact about its popularity or its ease of implementation. An API is an important attribute because it allows LLMs to act as players not by directly generating actions, but by producing programs that act as policies. Improvements in the code generation abilities of LLMs have allowed them to write small programs that can produce actions given game states without further intervention from the model. For instance, the VOYAGER system [38] leverages the code generation abilities of GPT-4 to play Minecraft (Mojang Studios, 2011) by interacting with the popular Mineflayer API. Using a sophisticated chain of prompts, VOYAGER generates blocks of code that leverage calls to the API in order to execute high-level “skills” (e.g. “Attack nearest zombie”) that are automatically converted into low-level game inputs (e.g. mouse movements and key presses). GPT-4 is also used as a high-level goal generator and planner, which in turn informs the code generation. This approach proved to be very successful, with VOYAGER being the first automated system to complete a variety of in-game Minecraft challenges. The results are impressive and indicate that generating action-producing programs may be a more efficient way to leverage latent LLM knowledge than direct action sampling. However, VOYAGER does benefit substantially from the availability of a robust API and vast amounts of internet discussions around Minecraft. As with the analysis of ChatGPT on Zork, the ability of this approach to generalize to less popular or entirely unseen games remains to be seen.
@@ 3.2 Non-Player Characters
Non-player characters (NPCs) are agents which exist in virtual game worlds but whose actions are not directly controlled by the players. NPCs exist to enrich the player’s experience and deepen immersion by adding to the world’s ambiance and making it more believable [40]. NPCs may serve as pets, allies, enemies [^2] We use enemies and allies here for in-game agents with different skills and ways to affect the world than e.g. game playing AI such as opponents in Chess., merchants, quest givers, or bystanders. Therefore, they have different agency even from AI-controlled players, and their goal is never to win. This makes designing AI for NPCs interesting [7], and LLMs are uniquely advantaged in this task since they can “understand” game-world settings and adapt their responses accordingly. It has been shown that LLMs are able to role-play through different scenarios [41, 42], thereby highlighting their potential to provide a more flexible and apt tool to emulate human behavior. We identify two ways in which LLMs can control NPCs: (a) through their dialogue, and (b) through their behavior. Behavior relates to in-game action selection, discussed in Section 3.1; however, we note that the heuristics and goals of such behavior is different than an AI player trying to win the game.
@@
LLMs are naturally suited for natural language conversation, and as NPC dialogue systems they can generate dynamic and contextually appropriate responses based on player input [43, 44, 45, 46]. This makes interactions with NPCs more engaging and realistic, reduces repetitive discourse and provides a more explorative experience within the game [47]. LLMs can engage the players in the gameworld’s narrative as foreground NPCs, background NPCs, or narrator NPCs. We discuss narrator LLMs as commentators in Section 3.4 whereas we cover the other two NPC types here. Foreground NPCs form part of the overarching narrative of the game, or one of its sub-narratives. They may be enemies, allies, information-givers, quest-givers, or item-providers. Their dialogue is heavily constrained by the scope of the narrative, their role within it, and the player actions. Foreground NPCs’ text generation process via LLMs must consider the overall context of the game and the interaction with the player, and keep track of events transpiring in the playthrough. This raises concerns regarding the memory capacity of LLMs, as well as the impact of possible hallucinations, i.e. plausible but false statements [48]. We revisit these limitations in Section 5. The purpose of background NPCs is to make the environment more believable and act independently of the players [47]. Such NPCs’ presence is purely decorative and their dialogue is essentially small talk. Thus their dialogue generation is less constrained, perhaps bound only by the identity of the speaker and their background. That said, their believability hinges on their ability to maintain the illusion that they have their own agency in the world and can interact with it [49]. Park et al. [43] explored NPC interactions within limited environments, where a number of LLM-based agents simulated social interactions in a sandbox environment. Within the constraints of their environment and the social affordances, the agents behaved in a believable manner, following their goals, planning for new ones, and even recalling past events as they interacted.
@@
Other studies have shown that multiple LLM-based agents are able to follow game rules and engage in game playing [50, 51], with different models consistently exhibiting their own aptitudes and weaknesses when applied to specific roles. This ability to interact within constraints is useful to instill believable behaviors in foreground and background NPCs, grounding their actions and dialogue within the rules of the game environment. Other work has focused more on the conversational and story-writing abilities of LLMs, such as the creation of dialogue between multiple characters, each having their unique personality, whilst following a consistent plot. One such example is the use of LLMs to generate a South Park (Comedy Central, 1997) episode [52] with multiple characters within a well-known setting. There are limitations to this approach, primarily that LLMs perform something like a theatrical improvisation, rather than acting as an actor studying a part [41]. Through this unconstrained process, the LLM is prone to hallucinations which do not fit the desired scenario. This volatility can be mitigated by providing the LLM not only with the conversation history but also with the current state of the environment, such as the items within it and their affordances, as well as the other characters and their corresponding actions. Urbanek et al. [53] used a configurable multi-user dungeon text-based environment in a fantasy setting to allow for both human and language model-based players. The latter were able to use an updated game state, including descriptions of local environments, objects, and characters, to take better actions and engage in coherent dialogue. This approach may also be extended to other scenarios or to cover the use of LLMs as active or interactive narrators. Ubisoft showcased LLM-based NPCs in their Neo NPC demo [46], where the player can freely converse with the in-game characters. Each NPC was given a carefully hand-crafted persona, and all NPCs could respond within the constraints of the game narrative and their prescribed personality whilst generating realistic responses. The players were able to engage in game-specific activities with these NPCs, such as planning a heist, or even attempt an unrelated course of dialogue altogether. Care was taken to minimize toxicity and social biases inherent in the LLMs through the prompt defining each NPC, which also had a bearing on how the latter would react to any player’s offensive or unruly discourse.